OpenStudy (anonymous):

find where the slope of the curve is defined: x^2y-xy^2=4

OpenStudy (anonymous):

does this mean "find the derivative, then check to see where the denominator is zero"?

OpenStudy (anonymous):

start with \[2xy+x^2y'-y^2-2xyy'=0\] solve for \(y'\) you will get a fraction then see for which \(x\) and \(y\) the denominator will be zero

OpenStudy (anonymous):

if that is not clear or if you need more help let me know

OpenStudy (anonymous):

sorrry I dont get it. Can you please explain the steps please?

OpenStudy (anonymous):

sure this is an implicit differentiation problem do you know how to find \(y'\) if \(x^2y-xy^2=4\) ?

OpenStudy (anonymous):

No I dont..

OpenStudy (anonymous):

ok then that is where we need to start you are thinking in this equation, that represents some curve, that \(y\) is a function of \(x\) even though we cannot solve for \(y\) in other words we are just imagining \(y\) as a function of \(x\) i.e \(y=f(x)\)

OpenStudy (anonymous):

now we can think of x^2y-xy^2=4 \[x^2y-xy^2=4\] instead as \[x^2f(x)-xf^2(x)=4\] and take the derivative of both sides the derivative of \(4\) is 0, so we will get a 0 on the right. the derivative on the left requires both the product rule and the chain rule so when we take the derivative of \(f^2(x)\) we will get \(2f(x)f'(x)\) by the chain rule

OpenStudy (anonymous):

using the product rule we get that the derivative is \[2xf(x)+x^2f'(x)-f^2(x)-2xf(x)f'(x)=0\] of course this is just an explanation, it is a lot easier to write \[2xy+x^2y'-y^2-2xy'=0\] let me know which if any step was confusing

OpenStudy (anonymous):

Okay I understand this part now.. but what do I do next?

OpenStudy (anonymous):

now that we have \[2xy+x^2y'-y^2-2xy'=0\] we need to solve for \(y'\)

OpenStudy (anonymous):

not too much here just some algebra \[x^2y'-2xy'=y^2-2xy\] \[(x^2-2x)y'=y^2-2xy\] \[y'=\frac{y^2-2xy}{x^2-2x}\] check my algebra

OpenStudy (anonymous):

and this will be undefined if the denominator is zero, so you last job, assuming that my answer is correct, is to solve \[x^2-2x=0\] which you can probably do with your eyeballs

OpenStudy (anonymous):

Thank you so much! I get it now. :)

OpenStudy (anonymous):

oh crap my algebra is wrong!!!

OpenStudy (anonymous):

hold on lets get it right, i made an algebra mistake

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

we start with \[2xy+x^2y'-y^2-2xyy'=0\] after taking the derivative using implicit differentiation, in this case in the form of the product and chain rule now we solve for \(y'\) carefully \[2xy+x^2y'-y^2-2xyy'=0\] \[x^2y'-2xyy'=y^2-2xy\] \[(x^2-2xy)y'=y^2-2xy\] \[y'=\frac{y^2-2xy}{x^2-2xy}\] that looks better

OpenStudy (anonymous):

now to solve \[x^2-2xy=0\] \[x(x-2y)=0\] so \(x=0\) or \(x=2y\) will make this undefined

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

you can ignore \(x=0\) because it is not in the domain of this expression

OpenStudy (anonymous):

if you replace \(x\) by \(0\) you get \(0=4\) in the original equation, so \(x=0\) is not on the curve anyway it is the other part that is your answer

OpenStudy (anonymous):

and of course you are still not done, because now that you have \(x=2y\) you have to solve in the original equation to see exactly what \(x\) and \(y\) are

OpenStudy (anonymous):

that is you have to solve \[x^2y-xy^2=4\] if \(x=2y\) you get \[4y^3-2y^3=4\] \[2y^3=4\] \[y^3=2\] \[y=\sqrt[3]{2}\]

OpenStudy (anonymous):

and you are STILL not done because you also have to solve for \(x\) but i will let you do that on your own. should be easy since you already have \(y\)

OpenStudy (anonymous):

Okay . Thank you so much! :)